OpenStudy (swagster23780):

What is the solution to the proportion? 9/x = 4/12 A.27 B.17 C.3 D.13

rishavraj (rishavraj):

wht do u think??

OpenStudy (swagster23780):

C

OpenStudy (theprofessor35):

Just cross multiply. Allow me to explain. :)

OpenStudy (theprofessor35):

9/x = 4/12 9*12 = 4x 4x = 108

OpenStudy (theprofessor35):

solve for x

OpenStudy (theprofessor35):

x = 27 Therefore, the missing value would be 27

OpenStudy (theprofessor35):

Your answer is 27. I hope I helped! xD